OpenStudy (jakyfraze):

What is the cube of 8? A. 512 B. 24 C. 64 D. 2

OpenStudy (jakyfraze):

can someone please explain:(

Nnesha (nnesha):

square = 2nd power cube =3rd power

Nnesha (nnesha):

and the power tells you how many times you should multiply the base example square of 3^2 means 3 times 3

OpenStudy (anonymous):

8*8*8 = 512

Nnesha (nnesha):

i think he can multiply. so let hiim do it #toomuchenabling

OpenStudy (jakyfraze):

So like 8*8*8?

Nnesha (nnesha):

how would you write it as cube power ?

Nnesha (nnesha):

cube of 8 =8 to the what power ?

OpenStudy (jakyfraze):

To the 3? I'm not really sure

Nnesha (nnesha):

yes cube =3rd power so 8^3 like i said exponent tells how many times u should multiply the base it's 3 then 8 times 8 times 8 is correct

OpenStudy (jakyfraze):

Please explain

Nnesha (nnesha):

here is an example 3^4 = 3 times 3 times 3 times 3 so 2^3 = 2 times 2 times 2 same way 8^3 = 8 times 8 times 8

OpenStudy (jakyfraze):

Oh my god thank you, I thought i got it wrong or said something stupid. Thank thank :)

Nnesha (nnesha):

np

OpenStudy (jakyfraze):

Do the answer is 512

OpenStudy (jakyfraze):

I'm your fan

Nnesha (nnesha):

yes thats correct

OpenStudy (jakyfraze):

thank you!!!

OpenStudy (jakyfraze):

i have another question, can i ask you please?

Nnesha (nnesha):

sure i'll try

OpenStudy (jakyfraze):

i want to close this so i can give you anther medal :)

OpenStudy (jakyfraze):

I'm going to open a new one after i close this one